OpenStudy (anonymous):

What are the similarities and differences between solving the equation, -3x + 2 = 11 and the inequality, -3x + 2 < 11? (Site 1) someone hurry im takin a SAT and i need the answer.

OpenStudy (anonymous):

-3x + 2 = 11 is an equation and it has only one solution. i.e. -3x= 11-2 i.e. -3x=9 x= 9/(-3) x=-3 Whereas -3x + 2 < 11 is an inequality and its solution will be -3x + 2 < 11 i.e. -3x < 11-2 i.e. -3x<9 i.e. x<9/(-3) i.e. x<-3 so it is obvious from th above solution that the solution of this inequality will all the numbers which are less than -3 i.e. x={.....-6, -5, -4}

OpenStudy (anonymous):

So an inequality has infinitely many solutions.

OpenStudy (anonymous):

@shonedell

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

How do you solve the inequality 12 < x + 5 so that the variable "x" is on the left side? (Site 2) help me with this one @dpasingh

OpenStudy (anonymous):

Ok wait

OpenStudy (anonymous):

12 < x + 5 i.e. 12-5<x i.e. 7<x i.e. x>7

OpenStudy (anonymous):

@shonedell

OpenStudy (anonymous):

in words

OpenStudy (anonymous):

@dpasingh